Equipment Status
Dome Camera: All Ok
Power Supply: Ok
Ring Switch: Ok
Findings: After raining water found in panel.
Suggestion: Solution for water presence.
Equipment Status
Dome Camera: All Ok
Power Supply: Ok
Ring Switch: Ok
Findings: After raining water found in panel.
Suggestion: Solution for water presence.
DTW-212 & 222 CCTV Maintenance Done Connector Change Now DTW-212 & 222 10 No’s CCTV Operation in NAV
IRP-1 CCTV monitor HDMI cable plug & power up Done
Daily Work Plan October 2022đ SL/NO Date PROBLEMS TYPE LOCATION REMARK 1 01.10.2022 DTW-105 CCTV checking maintenance. COR CCTV Team ICE: â PAPPU, Parvez,…
Optical Fiber Network Status October 2022 SL NO. Location In Between DTW Near Pole No Lat/Long Affected Location Remarks 01 Baherchor to Bottoli…
āĻĢā§āϞā§āĻ āϏā§āĻāĻ āĻāĻžāĻĻāĻžāϰ āϏāĻžāĻĨā§ āĻāĻāĻā§ āϝāĻžāĻāϝāĻŧāĻžāϰ āĻāĻžāϰāĻŖā§ āϞā§-āϞ⧠āĻāύā§āĻĄāĻŋāĻļāύ āĻĻā§āĻāĻžāϝāĻŧāĨ¤ āĻĒāϰ⧠āĻāĻ āϏāĻŽāϏā§āϝāĻž āĻāĻŽāϰāĻž āĻāĻāĻĄā§āύā§āĻāĻŋāĻĢāĻžāĻāĻĄ āĻāϰ⧠āϏāĻŽāĻžāϧāĻžāύ āĻāϰ⧠āĻĻāĻŋāϝāĻŧā§āĻāĻŋāĨ¤ āĻŦāϰā§āϤāĻŽāĻžāύ⧠āĻ āĻŋāĻ āĻāĻā§āĨ¤
One Comment